

********************************************************************
UARS WEEKLY STATUS REPORT - 02 JANUARY 1998
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems.
The half yearly calibration of the responsivity of the SUSIM working
channel have begun.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 355 CD-ROM requests as of
2-January-98. There were no new orders received or shipped this week.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
15 days PEM (3.32) Prod (1350R-725C=625L)
1 day CLAES (7.0) Eng (17R-1F=17L)
The CDHF will be implementing new login banners and issuing Rules of
Behavior for users. This activity is based on updated NASA and Goddard
security requirements.
Pending software upgrades: IDL 5.0, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 19 December 1997 (Orbit 34283, GMT Day 354) through 26
December 1997 (Orbit 34386, GMT Day 360).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 356/0535 34316 On
WINDII 357/0642 34331 On
WINDII 357/1002 34334 Off
WINDII 357/1600 34337 On
PEM 358/0026 34342 AXIS on
The beta angle reached a minimum angle of 0.0 degrees on 17 December 1997 (DOY
351). The beta angle is now increasing toward a maximum angle of 34.0 degrees
on 01 January 1998 (DOY 001).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.6 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.3 degrees C. Additional battery performance data for each battery on
20 December 1997 and 26 December 1997 are provided in Appendix D.
The clock error ranged from -11.4 to +10.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT 03
% HRDI 01
% MLS 01
% PEM 01
% WINDII 03
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT 02
TRANSPONDER FREQUENCIES:
Transponder A: 97/343 - 2287.497007
Transponder B: 97/344 - 2287.499121
GROUND SYSTEM OPERATION
Two (2) operational support problem(s) occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 97/352 show the observatory data loss to be 20 hours, 41
minutes, 12.916 seconds (no change). This is a 0.0378 percent data loss which
equals a 99.9622 percent data capture for the mission.
The FOT is still working out the details of participating in the IP return-link
telemetry test that the network is conducting.
No testing of the new CMS software release (R7B1V2) occurred this week.
UARS Test and Training Simulator (UTTS) testing of Q-channel data flow was
conducted on 22 and 23 December 1997. No successful OBC dumps were collected.
Several hardware changes were made by ATSC along the connection between the
UTTS VAX and the NASCOM lines. It is not certain what the cause of the Q-line
data flow problem is. Further testing is scheduled.
All On-Board Computer Software Tools Facility (OBCSTF) terminals and printer
were successfully connected and functioned properly on 23 December 1997. One
of two terminals and the laser printer had been experiencing communications
line problems with the VAX computer located in Building-14 after the move back
to the consolidated mission operations center.
Operational processing of level-0 telemetry analysis files on the UMPGAL
computer (new MPG Alpha machine) continues in parallel with the CDHF computer
system. Level-0 file transfers between CDHF and UMPGAL failed completely on 22
December 1997, preventing completion of level-0 processing on UMPGAL for that
day. File transfer problems continue to occur with a frequency of about two
failures per week. Unless this problem is rectified, level-0 processing on
UMPGAL will not prove to be practical. Parallel processing will continue
indefinitely, but the CDHF system will be the primary source for FOT telemetry
analysis products until the cause of the file transfer problem is isolated and
corrected. Until then, operational use of UMPGAL for processing will be
considered unreliable.
OTHER SIGNIFICANT EVENTS
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current status
of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 16 January 1998 at
016/13:25:14 UTC.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust will
probably be required in mid-1998.
FACILITIES
no change
PERSONNEL CHANGES
no change
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 December 1997
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
34284 354/0258 97-631 OBC Clk Adj fr 0.00 to 20.6 msec/day
34303 355/0910 97-632 OBC Clk Adj fr 20.6 to 0.00 msec/day
34316 356/0535 97-633 MLS Power ON
34331 357/0642 97-634 WINDII Power ON
34334 357/1002 97-635 WINDII Power OFF
34337 357/1600 97-636 HRDI Heater Delay to 100 Amp-min
34337 357/1600 97-637 WINDII Power ON
34342 358/0026 97-638 PEM Power ON AXIS
34368 359/1724 97-639 OBC/CLOCK Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 December 1997
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIEUTEMP Y-LO WINDII EU Temperature 34283-327
HRPSEQBIMON Y-HI HRDI Quiet Bus Current Monitor 34325,64,68
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
(none)
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 December 1997
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
34284 354/02:58 (20 Dec 97) from 00.0 to 20.6 msec/day
34303 355/09:10 (21 Dec 97) from 20.6 to 00.0 msec/day
34368 359/17:24 (25 Dec 97) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 December 1997
20 Dec 1997 - GMT Day 354 Beta = 8.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.7 0.000 0.0
2 +39.2/ -22.4 +5.47/+4.51 +32.0/-13.6 25.6 1.024 21.6
3 +112.0/ +0.0 +3.87/+2.92 +30.4/-15.2 25.6 1.021 20.9
26 Dec 1997 - GMT Day 360 Beta = 26.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.42/-3.68 +0.0/ +0.0 18.7 0.000 0.0
2 +44.8/ -5.6 +7.41/+5.79 +32.4/-16.0 24.6 1.023 23.6
3 +44.8/ -11.2 +6.11/+4.51 +31.6/-16.8 24.6 1.024 22.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 December 1997
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-124 FOT 34310 355/2028 14170 Generic Late Acquisition
97-124 FOT 34350 358/1238 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-126 FOT 34104 342/0231 TAC CRASH
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOS
------- ------ ----- ---- --- -------------
None this report period